OpenStudy (colors):

A new machine is offered at $12000 but can be acquired through installment at interest rate of 3.5% per month, compunded monthly. How much will u oay for intallment evert month for two years?

OpenStudy (retireed):

I think you mean 3.5% annual interest rate compounded monthly. It makes no sense to say 3.5% per month, compounded monthly.

OpenStudy (colors):

Ok can you please help to det. The how much will u pay for intallment every month for two years? @retirEEd

OpenStudy (retireed):

I learned this in college, so I will treat you like a college student. That means I will explain on a higher level than I would to a high school student. There is a fairly simple equation to calculate the payment. It is basically a loan repayment calculation like a car payment, but in this question we are just buying a new machine for $12,000. A simple Google search for "Loan Payment Formula" will find something like this equation.... \[Payment = r(PV)/(1-(1+r)^{-n})\] PV is the present value, (the loan amount, the cost of the new machine) r is the rate per annual period (the interest rate in decimal form divided number of periods per year n is the total number of periods to pay back the loan For your question; For your problem: PV = 12000 r = (0.035/12) = 0.002917 n = 24 be sure to use -24 like shown in the equation If you enter the information correctly your payment answer will be between $500 and $550